# Using cognitive models to reveal value trade-offs in language models

### v3

@@ 6 Discussion
While our approach offers several advantages, we also recognize the limitations of the cognitive models at the center of it. Cognitive models are often bespoke to the target domain they are crafted for, and so do not easily generalize to the open-ended nature of natural language use. Exploring how to use LLMs to map open-ended natural language data to the low-dimensional, interpretable feature space required for applying cognitive models (e.g. Jian & N, 2024; Qiu et al., 2025) will help to expand the settings we study with such models. Our initial attempts to do so revealed technical challenges at a number of points, first in getting smaller models to generate an alternative utterance set that captures the full semantic range of the true states we model, and later when trying to use the continuation probabilities of these generated utterances within the RSA model (c.f. Hu & Levy, 2023).
@@
Though the choices of values and goals used to construct the cognitive model in our work have been ecologically validated through human behavioral studies, they are certainly not the only goals that people entertain in communication, and further, might not be the particular set of goals that best describe LLM behaviors. Previous work has demonstrated that machine intelligence differs from our own (e.g. Schut et al., 2025), suggesting that human and machine conceptualizations of the world likely differ as well (Kim, 2022). One solution might be to develop new cognitive models of human-machine communication around neologisms that bridge human concepts and their machine counterparts to allow for a more precise understanding of LLMs as unique systems in their own right (cf. Hewitt et al., 2025).
Lastly, we hope that such theory-driven methods of interfacing with LLMs could also be of benefit to cognitive science. These large-scale learning systems could serve as a testbed for exploring how human social intelligence may have evolved, by allowing us to probe which trade-offs emerge naturally as a result of different architectures or training decisions, and which require explicit shaping(e.g. Bonawitz et al., 2019).
@@ B.5 Manipulations of communicative goal
Closed Models
Anthropic claude-3-5-sonnet-20241022 None
@@
Open Models
Qwen
(Qwen2.5-7B-Instruct)
HuggingFaceH4/ultrafeedback_binarized DPO
@@
[1pt/1pt]
Llama
(Llama-3.1-8B-Instruct)
HuggingFaceH4/ultrafeedback_binarized DPO
@@ C.2 Cognitive model
We generally follow the modeling assumptions described in Yoon et al. (2020), with one exception: where the original model assumes that negated expressions such as “not amazing” have more words and are thus slightly more costly for people to produce, we omit this additional cost and assume that each of the eight utterances are equally costly for an LLM.

### v4

@@ 6 Discussion
While our approach offers several advantages, we also recognize the limitations of the cognitive models at the center of it. Cognitive models are often bespoke to the target domain they are crafted for, and so do not easily generalize to the open-ended nature of natural language use. Exploring how to use LLMs to map open-ended natural language data to the low-dimensional, interpretable feature space required for applying cognitive models (e.g. Jian & N, 2024; Qiu et al., 2025) will help to expand the settings we study with such models. Our initial attempts to do so revealed technical challenges at a number of points (c.f. Tsvilodub et al., 2025), first in getting smaller models to generate an alternative utterance set that captures the full semantic range of the true states we model, and later when trying to use the continuation probabilities of these generated utterances within the RSA model (c.f. Hu & Levy, 2023).
@@
Though the choices of values and goals used to construct the cognitive model in our work have been ecologically validated through human behavioral studies, they are certainly not the only goals that people entertain in communication, and further, might not be the particular set of goals that best describe LLM behaviors. Previous work has demonstrated that machine intelligence differs from our own (e.g. Schut et al., 2025), suggesting that human and machine conceptualizations of the world likely differ as well (Kim, 2022). Developing new cognitive models of human-machine communication around neologisms that bridge human and machine-native concepts could allow for a more precise understanding of LLMs as unique system of their own (cf. Hewitt et al., 2025).
Lastly, we hope that such theory-driven methods of interfacing with LLMs could also be of benefit to cognitive science. These large-scale learning systems could serve as a testbed for exploring how human social intelligence may have evolved, by allowing us to probe which value trade-offs emerge naturally as a result of different architectures or training decisions, and which require explicit shaping (e.g. Bonawitz et al., 2019).
@@ B.5 Manipulations of communicative goal
Closed Models Anthropic claude-3-5-sonnet-20241022 None
@@
Open Models Qwen (Qwen2.5-7B-Instruct) HuggingFaceH4/ultrafeedback_binarized DPO
@@
[1pt/1pt] Llama (Llama-3.1-8B-Instruct) HuggingFaceH4/ultrafeedback_binarized DPO
@@ C.2 Cognitive model
We generally follow the modeling assumptions described in Yoon et al. (2020), with one exception: where the original model assumes that negated expressions such as “not amazing” have more words and are thus slightly more costly for people to produce, we omit this additional cost and assume that each of the eight utterances are equally costly for an LLM. Since LLMs were prompted to select from provided utterance options, we felt that it was more appropriate to not impose this notion of “cost” or “effort” of utterance production in the way it is conceived of in the cognitive model (the speaker’s effort in producing an utterance).